It is known that x + y = 10 and that x = z. Show that z + y = 10.

 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have

x+y=10 (i)

And, x=z (ii)

Applying Euclid’s axiom,

“if equals are added to equals, the wholes are equal”

We get,

From Eq. (i) and (ii)

x+y=z+y .(iii)

From Eqs. (i) and (iii)

z+y=10

Thanks